Twelve out of the 144 shoe city stores sell childern's clothing. You write the names of all of the stores locations on cards select a card at random. What is the probability of not picking a store location that sell children's clothing? justify.


please help!

Answer by stanbon(57387)

Twelve out of the 144 shoe city stores sell childern's clothing. You write the names of all of the stores locations on cards select a card at random. What is the probability of not picking a store location that sells children's clothing? justify.
---------
# of stores that do not sell children's clothing = 144-12 = 132
---
P(pick of store that does not sell children"s clothing) = 132/144 = 0.9167
==============
cheers,
Stan H.